OpenStudy (ayeedomo):

Please help with turning coordinates into function: (5,2),(6,5),(7,8),(8,11)

OpenStudy (ayeedomo):

@pooja195

OpenStudy (ayeedomo):

@caozeyuan

OpenStudy (caozeyuan):

what do you mean? coord into func?!

OpenStudy (ayeedomo):

I mean like the "y = __x + __"

OpenStudy (ayeedomo):

?

OpenStudy (caozeyuan):

ok but you don't need 4 pooints for that

OpenStudy (ayeedomo):

hmm

OpenStudy (caozeyuan):

ok, so these are on the same line

OpenStudy (caozeyuan):

so m =dy/dx, dy=5-2=3, dx=6-5=1, so m=3

OpenStudy (anonymous):